OpenStudy (kaley246):

Solve the following and round to the nearest tenth, if necessary. x2 + 5x - 2 = 0 (Points : 4) {-2, 5} {-5.4, 0.4} {-1, 0.5} No real solution

OpenStudy (kaley246):

@SolomonZelman

OpenStudy (solomonzelman):

This in the box, is the equation of the quadratic formula. \(\bbox[8pt, lime,border:8px solid black]{\LARGE \LARGE{x=~} \huge{ \frac{-\color{magenta}{b} \pm\sqrt{ \color{magenta}{b} ^2-4 \color{blue}{a} \color{red}{c}}}{2 \color{blue}{a}} } ~ }\) And this formula applies to a quadratic when the equation is as follows: \(\LARGE\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\)

OpenStudy (solomonzelman):

In your case, the equation is: \(\LARGE\color{black}{ \color{blue}{(1)} x^2+ \color{magenta}{(5)}x+ \color{red}{(-2)}=0 }\) and when you fill in the formula, you get: \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \LARGE{x=~} \huge{ \frac{-\color{magenta}{5} \pm\sqrt{ \color{magenta}{5} ^2-4 \cdot \color{blue}{(1)} \cdot\color{red}{(-2)}}}{2 \color{blue}{(1)}} }~ }}}\)

OpenStudy (solomonzelman):

@kaley246

OpenStudy (solomonzelman):

can you finish solving for x ?

OpenStudy (kaley246):

the answer is -5.4, 0.4

OpenStudy (kaley246):

right

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

correct!